# Optimal control transport of neutral atoms in optical tweezers at finite temperature

### v1

@@ III Conclusion
In the next steps, the assumptions on the noise can be further refined with experimental data toward an even more refined open-loop optimal control. The closed-loop optimal control also has now an input in terms of speed-limit potential expectation of how to use release-and-capture measurements as a figure of merit.
Code and data availability -- The code to perform the optimal control is available as QuOCS [22]. The code for transport is available upon reasonable request.
@@ Appendix A Simulation approach for transport
The numerical simulation for the atom transport uses a split-step method similar to the proposal in Ref. [18]. The main idea of the method is to consider the Hamiltonian of the system as a sum of the potential and kinetic parts, i.e.,
@@
The time-dependent Schrödinger equation $\mathrm{i}h\frac{\partial}{\partial t}\ket{\psi}=(H+V)\ket{\psi}$ can then be approximated via a Trotter decomposition of $K$ and $V$ for the unitary propagator of one time-step $dt$ as
$\exp\left(\frac{-\mathrm{i}dt}{h}(K+V)\right)$
$\approx\exp\left(\frac{-\mathrm{i}dt}{h}K\right)\exp\left(\frac{-\mathrm{i}dt}{h}V\right)\,.$ (8)
For the density matrix $\rho$, we evolve each state |$\ket{\psi_{i}}$ from Eq. (1) independently and recombine them with their weights $p_{i}$. The error in the approximation depends on the commutator $[K,V]$ and scales with the timestep as $\mathcal{O}(dt^{2})$; the error for the complete time evolution scales as $\mathcal{O}(dt)$ for the first order Trotter decomposition. The key mechanism are Fourier transformations implemented via the FFT library to have both components $K$ and $V$ in their diagonal eigenbasis. The exponential, now a diagonal matrix, becomes cheap and the FFT carries the major part of the workload [36]. The spatial grid is static over the evolution; no moving grid is required for the distances considered here.
@@
[…]
Figure 6: Convergence plots for various parameters. For every parameter, we show the convergence by comparing simulations at size $n$, $2n$, and $4n$ where $n$ is the default value of the parameter. The discretization of the spatial grid $dx$ shows a scaling of $\mathcal{O}(dx^{3})$, the one of the time grid scales as $\mathcal{O}(dt^{4})$. For the range, we increase the limits in the space $x_{\mathrm{max}}-x_{\mathrm{min}}$ at constant discretization $dx$. The cutoff $n_{c}$ is shown for $T=10\,\mu$K; we do not expect a linear scaling in log-log as the exponential distribution enters via the thermal state.
@@
State cutoff $n_{c}$ 2
For the convergence study, we take the setup from Tab. 1 to define the default size $n$ of the property to be converged. The specific simulations for each parameter lead to the convergence according to Fig. 6 and the infidelity $\mathcal{I}$ is defined as $\mathcal{I}=1-\mathcal{F}$. The cutoff $n_{c}$ is shown for $T=10\,\mu$K. We observe convergence with $dx$ following a scaling of $\mathcal{O}(dx^{3})$. For the time discretization the scaling is $\mathcal{O}(dt^{4})$. Neither the simulations concerning the grid extension nor in the cutoff show a polynomial convergence. For the latter, the behavior was expected as the exponential of the thermal state goes into the convergence. According to the results of the convergence study, we adopt the following parameters for the simulation: a time discretization of $dt=0.1\,\mu$s, a spatial discretization of $dx=0.002\,\mu$m, a grid extension of $x_{\max}-x_{\min}=10\,\mu$m, and a state cutoff of $n_{c}=8$.

### v2

@@ III Conclusion
The optimal control analysis can be extended to the simultaneous optimization of trap position, width, and depth. As a first future point, the assumptions on the noise can be further refined with experimental data toward an even more refined open-loop optimal control. The closed-loop optimal control also has now an input in terms of speed-limit potential expectation of how to use release-and-capture measurements as a figure of merit.
Code and data availability -- The simulations are based on the open-source software available in [38]. The data to reproduce the plots and all the figures are available at [39, 40].
@@ Appendix A Simulation approach for transport
The numerical simulation for the atom transport consists of computing the dynamics of the density matrix $\rho(t)$ under the application of the time-dependent potential $V(t)$. The time evolution of $\rho$ is performed by evolving each state |$\ket{\psi_{i}}$ from Eq. (1) independently and recombine them with their Boltzmann weights $p_{i}$. The statevector evolution uses a split-operator method similar to the proposal in Ref. [19]. The main idea of the method is to consider the Hamiltonian of the system as a sum of the kinetic and potential parts, i.e.,
@@
The time-dependent Schrödinger equation $\mathrm{i}\hbar\frac{\partial}{\partial t}\ket{\psi}=H\ket{\psi}$ can then be approximated via a Trotter decomposition of $K$ and $V$ for the unitary propagator of a small time-step $\text{d}t$ as
$\exp\quantity(\frac{-\mathrm{i}(K+V)\text{d}t}{\hbar})$
$\approx\quantity(\frac{-\mathrm{i}V\text{d}t}{2\hbar})\exp\quantity(\frac{-\mathrm{i}K\text{d}t}{\hbar})\exp\quantity(\frac{-\mathrm{i}V\text{d}t}{2\hbar})\,.$ (8)
The error in the approximation depends on the commutator $[K,V]$ and with this decomposition scales with the timestep as $\mathcal{O}(dt^{3})$. The key mechanism are Fourier transformations implemented via the FFT library to have both components $K$ and $V$ in their diagonal eigenbasis. The exponential, now a diagonal matrix, becomes cheap and the FFT carries the major part of the workload [41]. The spatial grid is static over the evolution; no moving grid is required for the distances considered here [19].
Figure 6: Convergence plots for various parameters. For every parameter, we show the convergence by comparing simulations at size $n$, $2n$, and $4n$ where $n$ is the default value of the parameter. The discretization of the spatial grid $dx$ shows a scaling of $\mathcal{O}(dx^{3})$, the one of the time grid scales as $\mathcal{O}(dt^{4})$. For the range, we increase the limits in the space $x_{\mathrm{max}}-x_{\mathrm{min}}$ at constant discretization $dx$. The cutoff $N_{s}$ is shown for $T=10\,\mu$K; we do not expect a linear scaling in log-log as the exponential distribution enters via the thermal state.
Figure 7: Figure of merit iterations. Optimal control iterations for $t_{\text{p}}=9\,\mu$s and $T=1\,\mu$K. The peaks represent the beginning of a new superiteration.
Figure 8: Evolution for different initial states with optimal control. (a) Ground state wavefunction before and after the evolution of duration $t_{p}=9\,\mu$s at $T=1\,\mu$K. (b) Expectation value of position $\langle x\rangle$ and momentum $\langle k\rangle$ for the ground state during the evolution. The inset contains the difference with the applied pulse $r(t)$ and shows the wavefunction oscillations. $\sigma_{x}$ and $\sigma_{k}$ represent the variance of position and momentum respectively. (c) Evolution of the second excited state. (d) The infidelity of the transport of $J_{\text{avg}}\sim 10^{-2}$ is not low enough to ensure the second excited state is not corrupted by the evolution.
@@
State cutoff $N_{s}$ 2
For the convergence study, we take the setup from Tab. 1 to define the default size $n$ of the property to be converged. The specific simulations for each parameter lead to the convergence according to Fig. 6 and the infidelity $\mathcal{I}$ is defined as $\mathcal{I}=1-\mathcal{F}$. The cutoff $N_{s}$ is shown for $T=10\,\mu$K. We observe convergence with $dx$ following a scaling of $\mathcal{O}(dx^{3})$. For the time discretization the scaling is $\mathcal{O}(dt^{4})$. Neither the simulations concerning the grid extension nor in the cutoff show a polynomial convergence. For the latter, the behavior was expected as the exponential of the thermal state goes into the convergence. According to the results of the convergence study, we adopt the following parameters for the simulation: a time discretization of $dt=0.1\,\mu$s, a spatial discretization of $dx=0.002\,\mu$m, a grid extension of $x_{\max}-x_{\min}=10\,\mu$m, and a state cutoff of $N_{s}=8$. Regarding the optimization, we fix a maximum number of 5000 iterations and 30 superiterations. Moreover, the number of parameters optimized at each superiteration is fixed as $N_{c}=t_{\text{p}}/10$ if $t_{\text{p}}\geq 20$ and $N_{c}=4$ for lower times. In the end, the total number of optimized parameters depends on how many superiterations are converged. On average, the optimization converges already at the first superiteration around 1000 as shown in Fig. 7.
For each statevector evolution, we store the wavefunction at each time step $t$ both in position and momentum space. We use this information to validate our results. For each initial state |$\ket{\psi_{i}}$, we analyze plots as the ones summarized in Fig. 8. Here, we report the example of an atom moved for $t_{\text{p}}=9\,\mu$s and at $T=1\,\mu$K via the optimal pulse $r_{\text{oc}}$. The figure of merit for this shuffling is $J_{\text{avg}}\sim 10^{-2}$. Specifically, Fig. 8(a) shows the probability density for the ground state $|\psi_{0}|^{2}$ before and after the evolution. We observe that the wavefunction remains localized in the trap. Moreover, since the transport is performed with a quite high fidelity, the expectation value of the position $\langle x\rangle$ has a small variance $\sigma_{x}$ and the momentum reaches zero again at the end of the evolution, without considerable oscillations as shown in Fig. 8(b). However, the evolution of the second excited state |$\ket{\psi_{2}}$ has a significantly larger standard deviation of the momentum. In Fig. 8(c-d), we see that the wavefunction |$\ket{\psi_{2}}$ at the end is distorted and delocalized. With this argument, we can qualitatively verify the chosen figure of merit and we can understand why a figure of merit $J_{\text{avg}}<10^{-2}$ is needed if we want to perform a high-fidelity transport even for the excited states.
Appendix B Transport distances analysis
In the main manuscript, we show the transport analysis for a fixed distance $d=3\,\mu$m. Here, we report the dependence of the quantum speed limit on the transport distance. In particular, we remember that the quantity $t_{\text{min}}$ is defined as the first minimum of the figure of merit which exhibits $J_{\mathrm{avg}}<10^{-2}$. First, we transport the atom via the piecewise quadratic pulse, and in Fig. 9(a), we observe that the behavior shows the same periodicity of $10\,\mu$s already highlighted in Fig. 2(a). For instance, with $r_{\text{pq}}$ we need the same time to transport the atom for $4\,\mu$m or $7\,\mu$m with a high fidelity. Instead, for the optimal control solution, the dependence becomes almost linear as shown in Fig. 9(b). Fluctuation of the linearity behavior may depend on the stochasticity of the optimization.
Figure 9: Quantum speed limit for different transport distances. (a) We report the minimum transport time with $J_{\mathrm{avg}}<10^{-2}$ for the piecewise quadratic pulse $r_{\text{pq}}$ at $T=1\,\mu$K. (b) With optimal control, the dependence of the quantum speed limit on the distance is almost linear.